# Theoretical analysis of photon detection mechanism in superconducting single-photon detectors

### v1

@@ III.3 Parameter region where the hotspot model is plausible
According to Ref. [24], where the authors assumed $\kappa=2$, they concluded that the simple hotspot expansion is a possible mechanism of photon detection. To compare it with the present work, we examine the dependence of $\kappa$ on $\Delta V$ and dynamics of the superconducting order parameter. We expect that the formation of the vortex–antivortex pairs becomes unstable as $\kappa$ decreases. As already shown in Fig. 4(b) for $\kappa=5$, we still observe the vortex–antivortex pair and the penetrating vortices and antivortices that have been observed for $\kappa=10$. We examine the case of $\kappa=1$, which is close to the transition ($\kappa_{c}=1/\sqrt{2}$) from type-II to type-I superconductor.
@@
Figure 8(a) shows the time evolution of $\Delta V$ for $\kappa=1$. We take $j_{\mathrm{bias}}/j_{0}=0.09$, which is slightly lower than the critical current. We observe the oscillatory behavior on $\Delta V$ especially for small $R_{\mathrm{init}}$. However, the profile is quite different from the previous one in Fig. 2 for $\kappa=10$: faster and slower oscillations coexist.
@@
One of the major difficulties in achieving photon detection using wide superconducting strips is the presence of intrinsic dark count as a result of the non-uniform distribution of the superconducting current in the strip. To examine the effect of the non-uniformity on the photon detection performance, we would like to ask how our result changes when the current flow becomes spatially more non-uniform. For this purpose, we examine the influence of $L_{y}$, the sample width, on the dynamics of the superconducting order parameter.
@@
Figure 10: Current density $j_{x}\left(x=L_{x}/2,y\right)$ before photon injection (dashed lines) and after the formation of the initial hotspot (solid lines) for $L_{x}\times L_{y}=40\xi_{0}\times 20\xi_{0}$ (red lines) and $L_{x}\times L_{y}=40\xi_{0}\times 40\xi_{0}$ (blue lines). Note that the integrated value of $j_{x}$ by $y$, $j_{\mathrm{bias}}$, is conserved. The current density after the hotspot formation is calculated at $t/t_{0}=150$. The initial hotspot radius is taken as $R_{\mathrm{init}}/\xi_{0}=3.0$ ($4.7$) for $L_{y}=20\xi_{0}$ ($40\xi_{0}$).
Figure 10 presents $j_{x}$ as a function of $y$ before and after photon injection. We consider a system with $L_{x}\times L_{y}=40\xi_{0}\times 40\xi_{0}$ which is twice the width of the other calculations (in this case we take $N_{y}=160$ with keeping linear mesh size $0.25\xi_{0}$). In the absence of the photon injection, we find that spatial non-uniformity has become more pronounced as $L_{y}$ increases. For $L_{y}=20\xi_{0}$, $j_{x}$ is almost parabolic as a function of $y$, but for $L_{y}=40\xi_{0}$, the parabolic nature is somehow deformed at both edges. With photon injection, $j_{x}$ is reduced at the location of the initial hotspot. For $L_{y}=40\xi_{0}$, $j_{x}$ is almost unchanged at the edge. On the other hand, for $L_{y}=20\xi_{0}$, $j_{x}$ outside the hotspot region increases after photon injection and tends to be spatially uniform. The result indicates that the spatial non-uniformity of $j_{x}$ outside the hotspot remains when $L_{y}$ is large.
@@
Based on these observations on the non-uniform current distribution for wider width cases, we examine the time-dependent voltage change $\Delta V$ and the dynamics of the superconducting order parameter for $L_{x}\times L_{y}=40\xi_{0}\times 40\xi_{0}$ to compare it with the previous results with $L_{x}\times L_{y}=40\xi_{0}\times 20\xi_{0}$. The voltage change is shown in Fig. 11. The threshold for successful photon detection is located at $R_{\mathrm{init}}/\xi_{0}=4.6$. The ratio between this threshold and the sample width $40\xi_{0}$ is $0.115$, and is smaller than the ratio $0.145$ in the case of $L_{y}=20\xi_{0}$ (the threshold is located at $R_{\mathrm{init}}/\xi_{0}=2.9$). Thus, the increase in the threshold is weaker than linear, even when we linearly increase the sample width. Near the threshold value, the oscillation of $\Delta V$ is still observed, although the signature of the oscillation is less clear than in the case of narrower width. This oscillation disappears as we increase $R_{\mathrm{init}}$, and this behavior is the same as that in the narrower-width case. Our observation, that the dependence of the threshold on $L_{y}$ is weaker than linear and the vortex–antivortex pair formation is a possible mechanism of photon detection near the threshold, is very helpful for the realization of SWSPD.
@@
The dynamics of the order parameter is presented in Fig. 12. We find that a vortex–antivortex pair is also formed in the initial hotspot area [see Figs. 12(a)(i) and 12(b)(i)], and then the antivortices penetrate the superconductor from the lower edge [see Figs. 12(a)(ii) and 12(b)(ii)]. After that, vortices and antivortices penetrate from the upper and lower edges repeatedly [see Figs. 12(a)(iii) and 12(b)(iii)]. The motion of the penetrating vortices and antivortices is more significant than that of the vortex–antivortex pair. As shown in Fig. 10, $j_{x}$ near both upper and lower edges of the sample after photon injection is almost independent of the sample width $L_{y}$, even though $j_{x}$ before photon injection is more non-uniform for $L_{y}=40\xi_{0}$ than for $L_{y}=20\xi_{0}$. This result suggests that the significance of the motion of the penetrating vortices and antivortices is not directly related to the non-uniformity of the current density in wider strips.
@@ IV Conclusion
Our numerical results and conclusions are somehow different from the previous works [24, 25]. Compared with the results in Ref. [24], even for large-$\kappa$ values, we confirm the usual hotspot mechanism, if $R_{\mathrm{init}}$ is large enough. In Ref. [24] with a small $\kappa$ value, the injection of sufficiently large photon energy generates vortex–antivortex pairs, and this is different from our results. Our model settings and parameter selection are much more systematic. Compared with the results in Ref. [25], the ratio between $R_{\mathrm{init}}$ and $\xi_{0}$ for photon detection due to the presence of vortex–antivortex pairs is different. In Ref. [25], the authors treated the case in which $R_{\mathrm{init}}$ is comparable to $\xi_{0}$, and did not focus on the threshold of $R_{\mathrm{init}}$. On the other hand, there exists the threshold value of $R_{\mathrm{init}}$ in our case, and their comparable setting does not show the finite amount of voltage change. In Ref. [25], the authors focused on the threshold bias current above which the superconducting state becomes unstable, and thus did not pay particular attention to the threshold of $R_{\mathrm{init}}$.

### v2

@@ III.3 Parameter region where the hotspot model is plausible
According to Ref. [Ota], where the authors assumed $\kappa=2$, they concluded that the simple hotspot expansion is a possible mechanism of photon detection. To compare it with the present work, we examine the dependence of $\kappa$ on $\Delta V$ and dynamics of the superconducting order parameter. We take $\kappa=1$, which is close to the transition ($\kappa_{c}=1/\sqrt{2}$) from type-II to type-I superconductor. We expect that the formation of the vortex–antivortex pairs becomes unstable as $\kappa$ decreases.
@@
Figure 8 shows the time evolution of $\Delta V$ for $\kappa=1$. We take $j_{\mathrm{bias}}/j_{0}=0.09$, which is slightly lower than the critical current. We observe the oscillatory behavior on $\Delta V$ especially for small $R_{\mathrm{init}}$. However, the profile is quite different from the previous one in Fig. 2 for $\kappa=10$: faster and slower oscillations coexist.
@@
A key difficulty in photon detection using wide superconducting strips is the intrinsic dark count arising from the non-uniform distribution of the transport current. It is therefore important to understand how our results change when the current flow becomes more spatially non-uniform in wider strips. To this end, we examine how the sample width $L_{y}$ affects the dynamics of the superconducting order parameter.
@@
Figure 10: Spatial profile of the supercurrent density $j_{s,x}(x=L_{x}/2,y)$ [upper panel (a)] and the order-parameter amplitude $\left|\tilde{\Delta}(x=L_{x}/2,y)\right|$ [lower panel (b)] before photon injection (dotted lines) and after the formation of the initial hotspot (dashed and solid lines) for $L_{x}\times L_{y}=40\xi_{0}\times 20\xi_{0}$ (red lines) and $40\xi_{0}\times 40\xi_{0}$ (blue lines). Since the normal-current contribution $j_{n,x}$ is small under the present conditions, the integral of $j_{s,x}$ over $y$ is approximately equal to $j_{\mathrm{bias}}L_{y}$. The supercurrent density and the order parameter after the hotspot formation are calculated at $t/t_{0}=150$ (dashed lines) and $t/t_{0}=200$ (solid lines). The initial hotspot radius is taken as $R_{\mathrm{init}}/\xi_{0}=3.0$ ($4.7$) for $L_{y}=20\xi_{0}$ ($40\xi_{0}$). We use $j_{\mathrm{bias}}/j_{0}=0.28$ and $\kappa=10$.
In Fig. 10(a), we compare the spatial profile of the supercurrent density, $j_{s,x}(L_{x}/2,y)$, before and after photon injection for a wider strip ($L_{y}=40\xi_{0}$) with those for $L_{y}=20\xi_{0}$. We note that all results presented so far were obtained for $L_{y}=20\xi_{0}$. For the wider strip, we keep the same mesh size $\delta=0.25\xi_{0}$, and in this geometry the critical current density lies between $j_{c}/j_{0}=0.285$ and $0.29$. For both strip widths, we set $j_{\mathrm{bias}}/j_{0}=0.28$ and $\kappa=10$. In the situations shown here, the normal-current contribution is small (although it is not explicitly plotted), so the integrated supercurrent $\int j_{s,x}\,dy$ is approximately equal to $j_{\mathrm{bias}}L_{y}$. Before photon injection [dotted lines in Fig. 10(a)], the screening effects—and hence the spatial non-uniformity of the current—become more pronounced as $L_{y}$ increases. For $L_{y}=20\xi_{0}$, the spatial profile of $j_{s,x}(L_{x}/2,y)$ remains close to $\cosh[(y-L_{y}/2)/\lambda]$ as expected from London theory. In contrast, for $L_{y}=40\xi_{0}$, an additional structure appears near the edges, reflecting the breakdown of the uniform-order-parameter assumption implicit in the London theory. For a large strip width $L_{y}$, the edge currents expected from the London theory become substantially larger than those that can be sustained by the self-consistent superconducting state. In practice, however, such large edge currents suppress the order-parameter amplitude through nonlinear effects [see blue dotted line in Fig. 10(b)], and this self-consistent suppression modifies the current distribution away from the $\cosh$ form near the edges. As a result, a distinct edge structure appears for $L_{y}=40\xi_{0}$.
After the photon injection [at $t/t_{0}=150$ and $t/t_{0}=200$, corresponding respectively to the dashed and solid lines in Fig. 10(a)], the supercurrent density inside the initial hotspot ($R_{\mathrm{init}}/\xi_{0}=3.0$ for $L_{y}=20\xi_{0}$ and 4.7 for $L_{y}=40\xi_{0}$) is reduced because the superconducting order parameter is suppressed by local heating, as shown in Fig. 10(b). In contrast, the current density surrounding the hotspot region increases, and is nearly saturated at $j_{s,x}(L_{x}/2,y)/j_{0}\lesssim 0.4$ for both strip widths. For $L_{y}=40\xi_{0}$, the additional structures of the supercurrent density near the edges, which are already present before photon injection as mentioned in the previous paragraph, evolve into local minima, whose values decrease with time. This behavior contrasts with the edge supercurrent density for $L_{y}=20\xi_{0}$. Such local minima correspond to the significant suppression of $\left|\tilde{\Delta}(L_{x}/2,y)\right|$ near the edges. We expect that the suppression affects the degree to which magnetic fluxes can penetrate.
@@
Based on these observations on the non-uniform current distribution for wider width cases, we examine the time-dependent voltage change $\Delta V$ and the dynamics of the superconducting order parameter for $L_{x}\times L_{y}=40\xi_{0}\times 40\xi_{0}$ to compare it with the previous results with $L_{x}\times L_{y}=40\xi_{0}\times 20\xi_{0}$. The voltage change is shown in Fig. 11. The threshold for successful photon detection is located at $R_{\mathrm{init}}/\xi_{0}=4.6$. The ratio between this threshold and the sample width $40\xi_{0}$ is $0.115$, which is smaller than the value $0.145$ for $L_{y}=20\xi_{0}$, with the corresponding threshold at $R_{\mathrm{init}}/\xi_{0}=2.9$. Thus, the increase in the threshold is weaker than linear, even when we linearly increase the sample width. Near the threshold value, the oscillation of $\Delta V$ is still observed, although the signature of the oscillation is less clear than in the case of narrower width. This oscillation disappears as we increase $R_{\mathrm{init}}$, and this behavior is the same as that in the narrower-width case. Our observation, that the dependence of the threshold on $L_{y}$ is weaker than linear and the vortex–antivortex pair formation is a possible mechanism of photon detection near the threshold, is very helpful for the realization of SWSPD.
@@
The dynamics of the order parameter is presented in Fig. 12. We find that a vortex–antivortex pair is also formed in the initial hotspot area [see Figs. 12(a)(i) and 12(b)(i)], and then the antivortices penetrate the superconductor from the lower edge [see Figs. 12(a)(ii) and 12(b)(ii)]. After that, vortices and antivortices penetrate from the upper and lower edges repeatedly [see Figs. 12(a)(iii) and 12(b)(iii)]. The motion of the penetrating vortices and antivortices is more significant than that of the vortex–antivortex pair. As shown in Fig. 10(b), $\left|\tilde{\Delta}(L_{x}/2,y)\right|$ is significantly suppressed near the edges for $L_{y}/\xi_{0}=40$, which faciliates the penetration of vortices and antivortices.
@@ IV Conclusion
Our numerical results and conclusions are somehow different from the previous works [Ota, Zotova]. Compared with the results in Ref. [Ota], even for large-$\kappa$ values, we confirm the usual hotspot mechanism, if $R_{\mathrm{init}}$ is large enough. In Ref. [Ota] with a small $\kappa$ value, the injection of sufficiently large photon energy generates vortex–antivortex pairs, and this is different from our results. Our model settings and parameter selection are much more systematic. Compared with the results in Ref. [Zotova], the ratio between $R_{\mathrm{init}}$ and $\xi_{0}$ for photon detection due to the presence of vortex–antivortex pairs is different. In Ref. [Zotova], the authors treated the case in which $R_{\mathrm{init}}$ is comparable to $\xi_{0}$, and did not focus on the threshold of $R_{\mathrm{init}}$. On the other hand, there exists the threshold value of $R_{\mathrm{init}}$ in our case, and their comparable setting does not show the finite amount of voltage change. In Ref. [Zotova], the authors focused on the threshold bias current above which the superconducting state becomes unstable, and thus did not pay particular attention to the threshold of $R_{\mathrm{init}}$.